# U-Motion: Learned Point Cloud Video Compression with U-Structured Motion Estimation

### v1

@@ 3.4 Spatial Predictive Coding
Besides inter-frame temporal redundancy, $f_{l}^{t}$ exhibits inter-layer spatial redundancy. Therefore, as shown in Figure 1, Spatial Predictive Coding module further removes the spatial redundancy by estimating an intra context $\bar{f}_{l}^{t}$, using the already decoded lower-scale latent $\hat{f}_{l-1}^{t}$.
@@
exercise_vox10 -26.41 -22.92 -30.47 -34.84
model_vox10 -40.35 -38.73 -34.78 -39.80
basketball_vox11 -3.93 -0.04 -25.05 -32.94
dancer_vox11 19.22 24.26 -2.76 -12.60
basketball_vox10 -26.17 -23.97 -31.46 -37.27
dancer_vox10 -12.36 -9.46 -17.65 -28.45
CTC-Overall -12.86 -9.35 -22.69 -30.04
Overall -14.99 -11.80 -23.31 -30.98
@@
The network detaches the redundancy in $f_{l}^{t}$ to generate residual $r_{l}^{t}$, then restores the reconstructed $\hat{f}_{l}^{t}$ from the decoded residual $\hat{r}_{l}^{t}$ using both inter context $\check{f}_{l}^{t}$ and intra context $\bar{f}_{l}^{t}$. An intuitive way of achieving this is through mathematical addition and subtraction: Given latent $f$ to be coded and the context $f_{c}$, generate residual $r=f-f_{c}$, and then recover the decompressed latent $\hat{f}=\hat{r}+f_{c}$, where $\hat{r}$ indicates the quantized $r$. To grant the network more flexibility, Li et al. [18] propose to use an encoder and decoder to detach and restore the context, respectively:
$r=\phi_{enc}(f|f_{c};\theta_{enc}),\hat{f}=\phi_{dec}(\hat{r}|f_{c};\theta_{dec}),$ (2)
where $\phi_{enc}$ and $\phi_{dec}$ are two neural networks. During experiments, we find that such modules are hard to train. Compared with mathematical subtraction that directly narrows ${r}$’s distribution, the distribution of ${r}$ generated by a encoder module is harder to control. Therefore, we propose to perform context detach with a residual connection:
$r=f-\phi_{enc}(f|f_{c};\theta_{enc}).$ (3)
The corresponding context restore operation is:
$\hat{f}=\hat{r}+\phi_{dec}(\hat{r}|f_{c};\theta_{dec}).$ (4)
The proposed context detach and restore operation grants the network with equivalent flexibility as Equation 2, but also provides the network with a ”shortcut” to narrow $r$’s distribution. Figure 2 shows the corresponding architecture of context detach and restore with residual connections. For motion embedding $e_{l}^{t}$, only one such module is used with context $e_{u,l-1}^{t}$. For latent $f_{l}^{t}$, a two-stage spatial-and-temporal context detach and restore module is employed to exploit the inter context $\check{f}_{l}^{t}$ and intra context $\bar{f}_{l}^{t}$.
@@ 3.5 Entropy Coding and Variable Rate Coding
The entropy coding module encodes the input $x$ ($x=e_{l}^{t}$ for motion compression and $r_{l}^{t}$ for latent compression) into a bitstream. We adopt the popular learnt compression approach, which uses both hyperprior and context features to estimate the probability distribution for quantized $x$ [3, 4]. See more detail in the supplement. For motion compression, the context $x_{c}=\left[e_{u,l-1}^{t},m_{l-1}^{t}\right]$. For latent compression, $x_{c}=\left[\check{f}_{l}^{t},\bar{f}_{l}^{t}\right]$. We adopt the global-and-local quantization scheme proposed in DCVC-FM [21], which generates the quantization stepsizes from a given rate control parameter $\lambda$. This enables a variety of bit rates with only one compression model.
@@
3.6 Geometry Compression
Our network can also be used for point cloud video geometry compression with slight modification. For geometry, the $l$-th layer point cloud $p_{l}^{t}=\left[G(p_{l}^{t}),O(p_{l}^{t})\right]$, where $O(p_{l}^{t})$ is an all-one vector indicating the occupancy. The same encoder structure is used to generate the latent $f_{l}^{t}$ and $\hat{f}_{l}^{t-1}$. To code $f_{l}^{t}$, because the actual geometry is not known yet at the decoder, we cannot generate the inter context $\check{f}_{l}^{t}$. Rather we first reconstruct $\hat{p}_{l}^{t}$ from $\bar{f}_{l}^{t}$, which is generated from $\hat{f}_{l-1}^{t}$ by sparse generative deconvolution with no target coordinates specified. The reconstruction is done by the APU module in Unicorn [31], which will estimate the probability that each voxel in $\bar{f}_{l}^{t}$ is occupied. For lossy compression, the top-N voxels in $\bar{f}_{l}^{t}$ with the highest occupancy probability will be kept to reconstruct $\hat{p}_{l}^{t}$; for lossless compression, arithmetic coder will be used to code $p_{l}^{t}$ losslessly using the estimated occupancy probability (thus $\hat{p}_{l}^{t}=p_{l}^{t}$). Based on the decoded $\hat{p}_{l}^{t}$ can we use the same U-Inter module shown in Fig. 1 to generate the temporal context $\check{f}_{l}^{t}$. Then using $\check{f}_{l}^{t}$ and spatial context $\bar{f}_{l}^{t}$, we code the current frame latent $f_{l}^{t}$ using the network architecture shown in Fig. 3. In other word, for geometry, ${f}_{l}^{t}$ is coded to help the reconstruction of $\hat{p}_{l+1}^{t}$ instead of $\hat{p}_{l}^{t}$.
@@ 4.1 Training and Testing Datasets
We follow the MPEG AI-3DG Common Test Condition (CTC), which specifies using the 8i Voxelized Full Bodies (8iVFB) [8] dataset for training, and the Owlii [17] dataset for testing. 8iVFB has a data precision of 10 bits and contains four sequences: soldier , longdress , redandblack and loot. Owlii contains four sequences in 11 bits precision: basketball_player, dancer, exercise, model. Following MPEG CTC’s requirement, we use 10-bit versions of model and exercise for testing provided by MPEG.
Figure 5: D1-PSNR performance comparison on lossy geometry compression among our method, Unicorn and D-DPCC
Figure 6: Ablation study on the contribution of motion flow.
@@ 4.2 Experimental Settings
where $R_{r,l}$ is the bit-rate of latent at layer $l$ and $R_{m,l}$ the bit-rate of motion. Note that $p_{l}^{t}$ is in YUV-space. We train only one model with $\lambda\in\left[256,18000\right]$. $\lambda$ is randomly picked from $\left[256,18000\right]$ for each iteration. We train our model with one NVIDIA A100-SXM4-80GB, and the overall training time is 4 days. During training, we use the lossless previous frame, i.e. $p^{t}$ as reference and only train one P-frame for each iteration. For geometry compression, the training details are in supplementary material.
@@
4.3 Results
Attribute Compression.
The R-D curves of attribute compression performance comparison is shown in Figure 4, with the corresponding BD-Rate detailed in Table 1. Please refer to the supplementary material for R-D curves with the YUV-PSNR. we assume the point cloud geometry is losslessly coded and not factored into the bit rate consumption. In addition, to remove the influence of I-frame, we use the original point cloud as I-frame, and do not consider the bits of I-frame when determining the bit rate (for U-Motion and baselines). We test the first 32 frames in each sequence. Each P-frame $p^{t}$ is coded using the reconstructed previous frame $\hat{p}^{t-1}$ as reference except for the first P-frame. Our method achieves significant gain against G-PCC-GesTM v3.0 and Unicorn on two 10-bit MPEG sequences exercise and model. Our method does not perform as well on 11-bit sequences basketball_player and dancer as on 10-bit sequences. We attribute this to the fact that the MPEG-specified training data are all 10-bit and have relatively smaller motion than these two 11 bits sequences, leading to a significant difference between training and testing data. When compared to the 10 bit versions of these two sequences, basketball_player_vox10 and dancer_vox10, our method achieve more significant gains. Figure 7 shows that U-Motion achieves better visual quality at a lower bit-rate compared to other methods (sequence model_vox10).
@@ 5 Ablation Study
5.1 Impact of Motion Estimation Methods
Figure 6 compares our method U-Motion with two variations, differing in motion estimation. For “zero motion”, we set the estimated motion to zero (the temporal context at each point is thus the interpolated feature of $K$ nearest neighbors of this point in the reference frame), and deduct the motion bit rate from the overall bit rate. Despite of the removal of the bits for coding motion features, the efficiency reduction of inter context leads to significant loss in R-D performance for the two 10-bit sequences. The performance loss is less pronounced for the two 11-bit sequences, likely due to the mismatch in the bit depth and motion characteristics of these sequences from the 10 bit sequences used in the training set. These results demonstrate the unequivocal benefits of explicit motion estimation. For “w/o bottom up”, we remove the bottom-up motion context by setting $e_{l}^{t}=e_{d,l}^{t}$. This leads to the increase in motion bits and reduces the overall R-D performance for all sequences and rate points. To assess the importance of the top-down connection in Figure 1, we trained a network without such a connection. We found that the network converges to spending no bits for coding motion, leading to similar results as zero-motion. We attribute this to the fact that 1) the lower-scale latent contains insufficient details for accurate motion estimation; 2) the inefficiency of compressing the high-scale motion without downsampling to lower-scale leads the network to allocate all bits for coding the latent. These results together reinforce the benefit of both top-down and bottom-up connections in the proposed U-structure for motion estimation under rate constraints.
@@ 5.2 Motion Visualization
Figure 8: Visualizing motion field from two viewing angles. The point cloud in the leftmost column moves from red to blue.
Figure 8 visualizes the motion estimated by our U-Inter module. The overlay of points of two frames on the left shows the movement between the two frames, where the hands and arms moving in opposite directions have the largest motion. This is reflected by the direction and magnitude plots of the estimated motion, demonstrating the effectiveness of the proposed U-Motion scheme.
Figure 9: Visualization of latent variable $f_{l}^{t}$, residual $r_{l}^{t}$ and the histogram depicting the distribution of $f_{l}^{t}$ and $r_{l}^{t}$.
5.3 Function of Context Detach
This section demonstrates the function of the context detach and restore. Figure 9 visualizes one variable in the latent $f_{l}^{t}$ before context detach and the residual $r_{l}^{t}$ after context detach. Compared with $r_{l}^{t}$, each region of $f_{l}^{t}$ shows relatively similar color distribution, indicating strong spatial dependencies. After contextual detachment, $r_{l}^{t}$ looks more like noise, indicating that most of the spatial redundancy is removed. We further show the distribution of one variable of $f_{l}^{t}$ and $r_{l}^{t}$. We can observe that $f_{l}^{t}$’s distribution is narrowed by contextual detachment.
@@ 5.4 Complexity
Table 3 compares the attribute compression complexity for our model, Unicorn and G-PCC-GesTM in terms of encoding and decoding times. U-Motion takes more time than Unicorn (about 25% more for 10 bit point cloud, about 85% for 11 bit) mainly due to the need for K-NN search for motion compensation. We use pytorch3d’s implementation of K-NN with $O(N^{2})$ complexity, leading to large amount of computation on point clouds with $0.7-3$ million points. However, the K-NN complexity can be greatly reduced if local K-NN applied. We compare the complexity of geometry compression in supplementary material.
@@ 6 Conclusion
We present U-Motion, a unified PCV codec achieving state-of-the-art performance for both attributes and geometry. The key innovation is to employ U-Structured motion estimation to capture motions of different scales under rate constraints. We also propose a novel context detach and restore module for efficient coding of motion and latent features. We show U-Motion’s applicability to both geometry and attribute, achieving significant gains over state-of-the-art methods. Our future work includes multi-frame training to alleviate error-propagation, reducing K-NN complexity for motion compensation and pretraining for U-Inter.
@@ 2.3 Results
Figure 13: Rate allocation between motion and latent over different layers corresponding to different rate points for basketball_player, frame index 00000016.

### v2

@@ 3.4 Spatial-Temporal Predictive Coding
Besides inter-frame temporal redundancy, $f_{l}^{t}$ exhibits inter-layer spatial redundancy. Therefore, as shown in Figure 3, Spatial-Temporal Predictive Coding module further removes the spatial redundancy by estimating an intra context $\bar{f}_{l}^{t}$, using the already decoded lower-scale latent $\hat{f}_{l-1}^{t}$.
@@
exercise_vox10 -29.69 -25.02 -33.75 -37.98
model_vox10 -43.62 -41.87 -38.32 -43.63
basketball_vox11 -4.05 1.89 -24.38 -32.50
dancer_vox11 20.92 28.94 0.04 -10.14
basketball_vox10 -26.72 -22.07 -32.40 -37.72
dancer_vox10 -16.43 -11.22 -22.05 -31.77
CTC-Overall -14.11 -9.84 -24.10 -31.06
Overall -16.58 -12.10 -25.14 -32.29
@@
Given latent representation $f$ to be coded and the context $f_{c}$, the network generates residual $r$ from the unpredictable part of $f$ given $f_{c}$. Inspired by the conditional coding in [19], we propose a context detach module with residual connection:
$r=f-\phi_{enc}(f|f_{c};\theta_{enc}),$ (3)
the corresponding context restore operation, which recovers $\hat{f}$ from the reconstructed residual $\hat{r}$ and context $f_{c}$, is expressed as:
$\hat{f}=\hat{r}+\phi_{dec}(\hat{r}|f_{c};\theta_{dec}),$ (4)
where $\phi_{enc}$ and $\phi_{dec}$ are two neural networks. The proposed context detach and restore operation grants the network with equivalent flexibility as conditional coding [19], but also provides the network with a “shortcut” to narrow $r$’s distribution. For motion $e_{l}^{t}$, context $f_{c}$ is $e_{u,l-1}^{t}$. For latent $f_{l}^{t}$, context $f_{c}$ is inter and intra context $\check{f}_{l}^{t}$ and $\bar{f}_{l}^{t}$.
@@ 3.5 Entropy Coding and Variable Rate Coding
The entropy coding module encodes the input $x$ ($x$ is the feature $r$ after the context detach module, $r=e_{l}^{t}$ for motion compression and $r=f_{l}^{t}$ for latent compression) into a bitstream. We adopt the popular feature compression approach, which uses both hyperprior and context features to estimate the probability distribution for quantized $x$ [3, 4]. See more detail in the supplement. For motion compression, the context $x_{c}=\left[e_{u,l-1}^{t},m_{l-1}^{t}\right]$. For latent compression, $x_{c}=\left[\check{f}_{l}^{t},\bar{f}_{l}^{t}\right]$. We adopt the global-and-local quantization scheme proposed in DCVC-FM [22], which generates the quantization stepsizes from a given rate control parameter $\lambda$. This enables a variety of bit rates with only one compression model.
Figure 5: D1-PSNR performance comparison on lossy geometry compression among our method, Unicorn and D-DPCC. The inconsistency of Unicorn and D-DPCC’s RD-curve compared with that in their original paper is due to different quantization methods used when downsampling 11-bit point clouds into 10-bit. Unicorn and D-DPCC [32, 33, 9] used $floor(\cdot)$ for quantization. Instead, We follow MPEG’s standard that uses $round(\cdot)$ for quantization. We have confirmed this with the authors of Unicorn and D-DPCC.
3.6 Geometry Compression
Our network can also be used for point cloud video geometry compression with slight modification. For geometry, the $l$-th layer point cloud $p_{l}^{t}=\left[G(p_{l}^{t}),O(p_{l}^{t})\right]$, where $O(p_{l}^{t})$ is an all-one vector indicating the occupancy. The same encoder structure is used to generate the latent $f_{l}^{t}$ and $\hat{f}_{l}^{t-1}$. To code $f_{l}^{t}$, because the actual geometry is not known yet at the decoder, we cannot generate the inter context $\check{f}_{l}^{t}$. Rather we first reconstruct $\hat{p}_{l}^{t}$ from $\bar{f}_{l}^{t}$, which is generated from $\hat{f}_{l-1}^{t}$ by sparse generative deconvolution with no target coordinates specified. The reconstruction is done by the APU module in Unicorn [32], which will estimate the probability that each voxel in $\bar{f}_{l}^{t}$ is occupied. For lossy compression, the top-N voxels in $\bar{f}_{l}^{t}$ with the highest occupancy probability will be kept to reconstruct $\hat{p}_{l}^{t}$; for lossless compression, arithmetic coder will be used to code $p_{l}^{t}$ losslessly using the estimated occupancy probability (thus $\hat{p}_{l}^{t}=p_{l}^{t}$). Based on the decoded $\hat{p}_{l}^{t}$, we use the same U-Inter module shown in Fig. 1 to generate the temporal context $\check{f}_{l}^{t}$. Then using $\check{f}_{l}^{t}$ and spatial context $\bar{f}_{l}^{t}$, we code the current frame latent $f_{l}^{t}$ using the network architecture shown in Fig. 3 (a). In other word, for geometry, ${f}_{l}^{t}$ is coded to help the reconstruction of $\hat{p}_{l+1}^{t}$ instead of $\hat{p}_{l}^{t}$.
@@ 4.1 Training and Testing Datasets
We follow the MPEG AI-3DG Common Test Condition (CTC) for dense dynamic point clouds, which specifies using the 8i Voxelized Full Bodies (8iVFB) [8] dataset for training, and the Owlii [18] dataset for testing. 8iVFB has a data precision of 10 bits and contains four sequences: soldier , longdress , redandblack and loot. Owlii contains four sequences in 11 bits precision: basketball_player, dancer, exercise, model. Following MPEG CTC’s requirement, we use 10-bit versions of model and exercise for testing provided by MPEG.
@@ 4.2 Experimental Settings
where $R_{r,l}$ is the bit-rate of latent at layer $l$ and $R_{m,l}$ the bit-rate of motion. Note that $p_{l}^{t}$ is in YUV-space. We train only one model with $\lambda\in\left[256,18000\right]$. $\lambda$ is randomly picked from $\left[256,18000\right]$ for each iteration. We use $5$ layers to construct our network, i.e. $K=4$. The highest scale $H$ equals to the precision of data (11/10 for 11/10-bit point clouds). We choose not to have a U-Inter module in the highest scale $H$ due to complexity issue, as the K-NN search in scale $H$’s motion compensation brings excessive computational cost. We use only the spatial context $\bar{f}_{H}^{t}$ updated from $H-1$ scale’s decoded feature $\hat{f}_{H-1}^{t}$ when coding scale $H$. We train our model with one NVIDIA A100-SXM4-80GB. The overall training time is 4 days. During training, we use the lossless previous frame, i.e. $p^{t}$ as reference and only train one P-frame for each iteration. For geometry compression, the training details are in supplementary material.
@@
4.3 Results
Attribute Compression.
The R-D curves of attribute compression performance comparison is shown in Figure 4, with the corresponding BD-Rate detailed in Table 1. Please refer to the supplementary material for R-D curves with the YUV-PSNR. We assume the point cloud geometry is losslessly coded and not factored into the bit rate consumption. In addition, to remove the influence of I-frame coding, we use the original point cloud as I-frame, and do not consider the bits of I-frame when determining the bit rate (for U-Motion and baselines). We test the first 32 frames in each sequence. The original first frame is used as the decoded I-frame, and each subsequent frame is coded as a P-frame using U-Motion. Our method achieves significant gain against G-PCC-GesTM v3.0 and Unicorn on two 10-bit MPEG sequences exercise and model. Our method does not perform as well on 11-bit sequences basketball_player and dancer as on 10-bit sequences. We attribute this to the fact that the MPEG-specified training data are all 10-bit and have relatively smaller motion than these two 11 bits sequences, leading to a significant difference between training and testing data. When compared to the 10 bit versions of these two sequences, basketball_player_vox10 and dancer_vox10, our method achieve more significant gains. Figure 6 shows that U-Motion achieves better visual quality at a lower bit-rate compared to other methods (sequence model_vox10).
@@ 4.3 Results
Figure 7: Ablation study on U-Inter’s different modules.
@@ 5 Ablation Study
5.1 Impact of U-Inter
Figure 7 compares our method U-Motion with four variations, differing in motion estimation and coding.
Gain Brought by Motion.
As the baseline codec Unicorn uses no motion information, we train a variation of the U-motion model with “zero motion”. The temporal context at each point is the interpolated feature of its $K$ nearest neighbors in the reference frame. Despite zero motion bits used, the temporal context obtained with zero motion leads to significant loss in R-D performance. These results demonstrate the unequivocal benefits of explicit motion estimation.
Figure 8: Group motion flow magnitude visualization. In the ”Combination” column, gray points belong to reference frame, whereas black points belong to current frame. Warmer color in the visualization of motion denotes larger movements.
Top-Down Motion Propagation.
For “w/o top-down”, we trained a model without the motion propagation block in Fig. 2 (i.e. $e_{l}^{t}=\tilde{e}_{l}^{t}$ in Fig. 2). Note that this blocks the motion information propagation from fine to coarse layers, which restricts the compression of motion information into its own scale. We found that only the lowest scale contains non-zero motion bitstream ($r_{e,H-K}^{t}$) with this variation, leading to overall significant performance degradation.
Bottom-Up Motion Predictive Coding.
In the “w/o bottom up” variation, we remove the bottom-up “Motion CTX Detach“ and “Motion CTX Restore“ in Fig. 2 for coding the motion information. This leads to increased motion bits and reduced R-D performance for all sequences and rate points.
Multi-scale Motion Compensation.
For “single-scale MC”, we only perform motion compensation (MC) at scale $H-2$ (8 for 10 bit point cloud) to generate the temporal context. For other scales, only spatial context is used. We choose scale $H-2$ for MC because this is the same scale as in previous point cloud video compression methods [9, 17]. Note that although we only do MC in scale $H-2$, we still do U-structured motion estimation, propagation and coding from scale $H-4$ to $H-2$ to accurately estimate the scale $H-2$ motion. We can observe significant gain brought by multi-scale MC due to the rich temporal context brought by multi-scale MC that strengthen the coding of each layer.
Group Motion.
For the “w/o group motion” variation, we change group motion to one-motion-per-scale (group size=1). We can witness a performance drop due to the decay of expressiveness of the motion flow, especially in sequences with larger movements (model_vox10).
@@ 5.2 Motion Visualization
Figure 8 visualizes the decoded group motion generated by our U-Inter module at different scales. Column 4-7 shows the motion (for group 0) from scale $H-4$ to scale $H-1$. We can observe that 1) finer-scale motion contains more details than coarser-scale motion; 2) finer-scale motion are more consistent and physically correct than coarser scale for areas with large motion. We attribute this to the fact that finer-scale features contains more high-frequency details for correct motion estimation. However, if we directly estimate the motion at scale $H-1$, we do not get accurate results because 1) the perceptive field of local convolution limits the estimation of large motions; 2) it’s inefficient in rate-distortion to compress scale $H-1$’s motion directly.
For different motion groups, we observe that the network tends to make different predictions among groups in some areas with large movements (red circle). The network uses different motion predictions in these areas to enhance the temporal context, leading to the performance gain compared with single-group motion estimation.
@@ 5.3 Complexity
Table 3 compares the attribute compression complexity for our model, Unicorn and G-PCC-GesTM. U-Motion takes more time than Unicorn (about 25% more for 10 bit point cloud, about 85% for 11 bit) mainly due to K-nearest-neighbor (K-NN) search for motion compensation. We use pytorch3d’s implementation of K-NN with $O(N^{2})$ complexity, leading to large amount of computation on point clouds with $0.7-3$ million points. We can observe in Table 3 that K-NN takes up large amount of the encoding/decoding time. However, the K-NN complexity can be greatly reduced if K-NN search is restricted to a local region. The geometry compression complexity is described in the supplementary material.
@@ 6 Conclusion
We present U-Motion, a PCV codec employing both U-structured temporal prediction and coarse-to-fine spatial prediction. The key innovation in U-Motion is its U-inter module, combining multi-scale motion estimation, fine-to-coarse motion propagation and coarse-to-fine predictive coding of motion, to enable effective multi-scale motion compensation under rate constraints. We show U-Motion’s applicability to geometry and attribute, achieving significant gains over state-of-the-art methods. Our future work includes multi-frame training to alleviate error-propagation and reducing K-NN complexity for motion compensation.
@@ 2.3 Results
Figure 12: Rate allocation between motion and latent over different layers corresponding to different rate points for model, frame index 00000150.